# Weakly Supervised Object Segmentation by Background Conditional Divergence

### v1

@@ Acknowledgments
Research at the University of Delaware was sponsored by the Department of the Navy, Office of Naval Research under ONR award number N00014-21-1-2300. This material is based upon work supported by the National Science Foundation under Award No. 2108841. This research was also supported in part through the use of DARWIN computing system: DARWIN – A Resource for Computational and Data-intensive Research at the University of Delaware and in the Delaware Region, which is supported by NSF under Grant Number: 1919839, Rudolf Eigenmann, Benjamin E. Bagozzi, Arthi Jayaraman, William Totten, and Cathy H. Wu, University of Delaware, 2021, URL.
@@ A.1 Generative-based Models
The work by Remez et al. (2018) is similar to ours in that it proposes a weakly-supervised approach that relies on a discriminator to assess the statistical similarity between synthetic images with counterfactual backgrounds and real images. Notably, their discriminator relies on a pretrained object classifier. Furthermore, the masking network input is based on the features from a pretrained object recognition network which outputs a bounding box for the foreground object. The necessity of using pretrained models and a network to generate a bounding box are the main limitations of this work.
@@
Bielski & Favaro (2019) proposed a GAN-based, unsupervised approach to segment foreground pixels. The GAN produces mask, foreground, and background images. The mask is used to alpha-blend the background and foreground back together. Trivial solutions for this work result when (i) the generated background and foreground are identical, so any arbitrary mask would still give a real image, (ii) Either the foreground or the background is ignored (i.e., all-foreground or all-background), which results in an either all-zero or all-one mask. To prevent the latter, the hinge loss on the average mask value is deployed. To prevent the former, a small random spatial shift of the composite image and mask are applied before alpha-blending to create a counterfactual shift in the resulting synthetic image. This shift penalizes having poor generated masks with identical foreground and background by creating a statistical divergence between shifted counterfactuals (i.e., the generated image) and real images, which the discriminator can detect. A drawback to this approach is the number of different loss functions (size, binary, gradient penalty, discriminator output norm) and hyper-parameters required to prevent degenerate solutions.
@@
The work by Sauer & Geiger (2021) introduced a Counterfactual Generative Network (CGN), leveraging the concept of learning Independent Mechanisms (IMs) to establish a causal structure for mask, foreground, and background. Each was modeled by fine-tuning a distinct pretrained BigGAN. To ensure the resulting composite image is real, they minimize the MSE loss between the output of pre-trained frozen BigGAN and the composition process. To ensure the generated mask is not trivial, hinge and entropy losses are deployed. For the background loss, the predicted saliency is minimized which leads to the foreground object’s disappearance. For background generation a pre-trained U2Net is used to detect the saliency. Overall, this approach heavily relies on pretrained models and necessitates the incorporation of specific losses to prevent degenerate or trivial solutions.
@@ A.2 Vision-Transforms-based Models
The work by Hamilton et al. (2022) proposed an unsupervised semantic segmentation approach that distills knowledge from pretrained feature extractors. A frozen backbone (e.g., DINO-ViT) is used to extract dense pixel-wise features from input images. A trainable segmentation network maps these features to lower-dimensional embeddings representing per-pixel segment assignments. During training, a mini-batch is constructed for each sample by including its K-nearest neighbors in the feature space. The model then samples a neighbor image and computes a cosine similarity matrix between the pixel features of the target and the neighbor. The segmentation output is encouraged to be similar for pixels with high feature similarity and dissimilar for pixels with low similarity
@@ A.3 Class-Activation-Map-based Models
Much work on weakly-supervised segmentation is based on class-activation maps (CAM) derived from pretrained convolutional object classification networks Kumar Singh & Jae Lee (2017); Wei et al. (2017); Zhang et al. (2018); Li et al. (2018); Hou et al. (2018); Jiang et al. (2019). We stress the fact that our model does not require labeled data (for instance, class information for composite/background images). Performance was initially limited as segmentation based on activation maps often focused on a few parts of the objects (Kumar Singh & Jae Lee, 2017; Wei et al., 2017). To mitigate this problem, Zhang et al. (2018) used a backbone fully-convolutional network to extract initial CAMs and then used two parallel-branches, each a CNN classification network, which took the initial CAM as an input. Parts of the input CAM of the first classifier are erased with the guidance of the object localization map from the second classifier. This encourages the first classifier to seek other parts of the image to do the classification. Thus, this results in learning a more coherent/complete feature map that can be used for object segmentation. Similarly, Li et al. (2018) proposed two networks, where one network uses the parts that are neglected by the first one for the classification task which results in expanding the activation map to other related parts. The work by Hou et al. (2018) introduced an approach for adversarial learning during supervised classification training which encourages the CAM to expand from covering only part of the target objects while penalizing it for covering background regions spuriously using the background prior (i.e., regions with low activations defined as background, regions with high activations defined as foreground and regions in-between defined as potential foreground) obtained from the initial map Following a similar aim, Jiang et al. (2019) observed that the attention maps produced by a classification network changes focus on different object parts throughout the learning. They saved copies of the model each epoch and then used them to produce CAMs. At the end, they fused all the resulting CAMs using a pixel-wise max operation, which enhanced the final CAM. All previously mentioned work in this paragraph depends on pretrained models.
@@ A.4 Other Work
Hung et al. (2019) proposed SCOPS, a self-supervised part segmentation framework that leveraged multiple auxiliary objectives. First, they used saliency maps extracted from a pre-trained classification network as soft supervision, encouraging the encoder’s part activation maps to align with semantically meaningful regions via a cosine similarity loss. Second, they introduced a geometric loss that encouraged all pixels assigned to a given part to be spatially concentrated around that part’s center; the number of parts was treated as a hyperparameter. Lastly, an equivariance loss ensured that part assignments remained consistent under random spatial transformations such as rotation and scaling, promoting robustness and part consistency.
The work by Savarese et al. (2021) proposed an unsupervised model-free method to segment foreground and background based on minimizing the mutual information between the segmented foreground and background. They used a variant of information theory, namely the coefficient of constraint that prevents trivial solution including an entropy in the denominator. By including assumptions about the conditional distributions (i.e., Laplace distribution for foreground/background given background/foreground), the loss is equivalent to maximizing the inpainting errors of both the masked foreground and the masked background. However, they added a new loss that penalizes the texture diversity within the masked region (i.e., the masked region should have a uniform texture)
@@ Appendix B Statistical Divergences
The family of Wasserstein distances is another class of divergences based on optimal transport theory (Villani et al., 2008). In particular, the Wasserstein-$1$ distance, a statistical divergence also known as the Earth mover’s distance, has been applied in generative modeling in the form of the Kantorovich dual (Arjovsky et al., 2017). Note that all of these divergences require adversarial optimization (MMD-GAN (Li et al., 2017) uses adversarial selection of the kernel). In contrast, the primal form of the Wasserstein-$p$ distance measures the divergence between the distributions of two random variables $X\sim\mathbb{P}_{X}$ and $Y\sim\mathbb{P}_{Y}$ in terms of the optimal transport plan
@@
[…]
@@
Given two samples/batches ^$\hat{\mathbb{P}}_{X}=\{(\frac{1}{N},\mathbf{x}_{i})\}_{i=1}^{N}$ and ^$\hat{\mathbb{P}}_{Y}=\{(\frac{1}{N},\mathbf{y}_{i})\}_{i=1}^{N}$, the computation of the sliced Wasserstein-$p$ distance is simply the mean of the absolute error to the $p$th power between paired samples after sorting: $\mathrm{SW}_{p}^{p}(\hat{\mathbb{P}_{X}},\hat{\mathbb{P}}_{Y};\mathbf{w})=\frac{1}{N}\sum_{j=1}^{N}(\mathbf{w}^{\top}\mathbf{x}_{\pi_{j}}-\mathbf{w}^{\top}\mathbf{y}_{\sigma_{j}})^{p}$, where the permutations $\pi$ and $\sigma$ ensure that $\mathbf{w}^{\top}\mathbf{x}_{\pi_{1}}\leq\mathbf{w}^{\top}\mathbf{x}_{\pi_{2}}\leq\cdots\leq\mathbf{w}^{\top}\mathbf{x}_{\pi_{N}}$ and $\mathbf{w}^{\top}\mathbf{y}_{\sigma_{1}}\leq\mathbf{w}^{\top}\mathbf{y}_{\sigma_{2}}\leq\cdots\leq\mathbf{w}^{\top}\mathbf{y}_{\sigma_{N}}$, respectively. Intuitively, $\mathbf{w}$ defines a subspace and the sorting ensures the shortest distances between the pairs in the subspace.
Alternatively, once can use energy-based Wasserstein (EBSW) (Nguyen & Ho, 2023) that depends on importance sampling to emphasize slices that distinguish the distributions:
@@
[…]
@@
Interestingly, an analogous approach can be used to select from a predefined set of kernel functions for MMD (Biggs et al., 2023), which is provides a rough approximation of the adversarial optimization of the kernel function in MMD-GAN (Li et al., 2017).
@@ C.1 Synthetic Data Description
C.1.1 Synthetic-Aperture Sonar Backgrounds with dSprites
This synthetic dataset is derived from real-world background-only synthetic-aperture sonar (SAS) images (Cobb & Zare, 2014), augmented by overlaying artificial objects to simulate relevant scenarios. This SAS dataset contains 129 $1001\times 1001$ images that are originally complex-valued, single (high-frequency) channel, taken from various seafloor texture types. For each sonar scene, we crop 200 $64\times 64$ images from the magnitude images. We normalize each crop by its mean and then clip any value larger than 16. Half of the resulting crops are used as composite images where we paste objects from dSprites (Matthey et al., 2017) dataset using equation 1. The data is divided into $70\%,20\%,\text{ and }10\%$ for training, validation, and testing.
@@ C.1.1 Synthetic-Aperture Sonar Backgrounds with dSprites
C.1.2 Textures with MNIST, FashionMNIST, and dSprites
@@
For the synthetic dataset, $140,000$ background images of size $64\times 64$ were generated by randomly patching the texture images. Half of the background images were used as background for the composite images. This resulted in $70,000$ images as foreground and background. Images in the foreground datasets were resized to $64\times 64$ and the true (binary) mask was used to blend it onto the background image using equation 1. Since MNIST and Fashion MNIST have no given true mask, we estimate the true mask by applying Otsu’s threshold on each image (Otsu, 1979). The data is divided into $70\%,20\%,\text{ and }10\%$ for training, validation, and testing.
@@
C.2 Implementation Results
@@
For SAS+dSprites and Textures+MNIST/FashionMNIST/dSprites, the surrogate data cluster labels for the composite images are computed by processing the corresponding background-only images through both the AE/CLE and Sinkhorn K-means. For the SAS images, we find that AE architecture is not able to reconstruct the images, so we use a symmetrical architecture that is composed of four blocks for the encoder and decoder, and each block is composed of convolution, ReLU, BatchNorm, with residual units: the 64-channel block has 1 and other blocks have 2 residual units each. The number of channels per block is $\{64,128,256,512\}$, with an additional 768-channel stage. As described above, a fully connected bottleneck layer of dimension $o=20$ is used to represent the latent space.
@@
The Textures+MNIST/FashionMNIST/dSprites images, have 10 ground truth clusters and we experiment with varying the number of clusters, choosing values from the set $\{5,10,15\}$
For SAS images, we choose cluster values from the set $\{3,4,5\}$. Initial experiments with 2 clusters consistently failed to separate visually distinct regions.
@@
For Textures+MNIST/FashionMNIST/dSprites and SAS+dSprites datasets, we run the training for $50$ and $100$ epochs, respectively. For the former, we choose the model with lowest validation loss $\mathcal{L}$ shown in equation 8, but for the latter we also include a quantile loss equation 15 with quantile, $q=0.8$, such that the maximum area of an object is 20% of the image size:
@@
$\mathcal{L}_{\text{quantile}}(\theta)$ $=\mathbb{E}_{X,R}\left[\sum_{ij\in\mathcal{Q}}\big{(}[\tilde{X}_{\theta}]_{ij}-R_{ij}\big{)}^{2}\right],\quad\mathcal{Q}$ =$=\left\{ij:M_{\theta}(X)_{ij}\leq\tau\right\},\quad\tau$ $=\text{quantile}\left(M_{\theta}(X),q\right).$ (15)
[…]
C.3 Results: Synthetic Object on Real Sonar Dataset
We present both quantitative and qualitative results for the sonar dataset that combines real-world seafloor imagery from Cobb & Zare (2014) with synthetic objects from the dSprites dataset. The seafloor contains six texture categories: sand ripple, rock, coral, sand, mud, and seagrass.
Figure 5 illustrates the segmentation performance, demonstrating that the proposed method generates high-quality segmentation masks. However, for this dataset, the method does not produce ‘sharp’ masks.
Figure 5: Results of the segmentation model for sonar images. From top to bottom: input images, estimated masks, true masks, and thresholded estimated masks at .5.
Table 5 provides the performance metric values for different numbers of background clusters. The results indicate that the best performance is achieved when using four clusters.
Table 5: Performance Metrics Across Different Number of Clusters for Sonar Data
Clusters Avg. Precision (%) IoC (%) AUC-ROC (%)
3 60 30 90
4 82 57 97
5 22 20 87
C.4 Results: MNIST, FasionMNIST, and dSprites on Textures
We provide both quantitative and qualitative results for three datasets: MNIST, Fashion MNIST, and dSprites. Initially, we trained the segmentation model using true clusters, defined by the texture label. Subsequently, we trained the segmentation model based on AE clustering. We begin by presenting quantitative results, followed by showcasing qualitative outcomes for the best-performing models on each dataset.
Figure 6: The percentage of Average Precision, AUCROC, IOC values (y-axis) for the four datasets, as indicated by the legend. The x-axis represents the chosen number of clusters for the K-means model based on AE/CLE features. The star symbols above the x-axis values, specifically at the position of 10, emphasize the Average Precision and AUCROC corresponding to the ground truth clusters.
We refer to the evaluation values corresponding to the model trained using the true clusters by the ground truth values. We show 8 testing images, the estimated masks, the true masks, and the thresholded estimated mask based on $0.5$ in Figure 7 using 10 clusters obtained from the AE+Kmeans. Figure 6 shows the results for the three datasets with respect to the number of clusters from K-means. The star shows the performance using the ground truth values (i.e., the true texture class). To qualitatively compare the performance of true clusters vs the estimated clusters in the latent space using K-means, we show the segmentation performance for the true clusters in Figure 9
The performance of the model on Fashion MNIST is high with all clusters performing relatively the same. The Fashion MNIST objects are relatively large with polygonal-like shapes and the masks are accurate as shown in Figure 7.
Figure 7: Results of the segmentation model for Fashion MNIST (left), MNIST (middle), and dSprites (right). From top to bottom: input images, estimated masks, true masks, and thresholded estimated masks at .5
The model performs similarly on MNIST when utilizing any synthetic clusters, except for a slight drop when using the 10 clusters. We notice, however, that the performance of the true clusters is worse compared to the synthetic one. As shown in Figure 9, the performance drops when using true clusters.
The model’s performance on the dSprites dataset is comparatively lower, which we primarily attribute to the small size of objects and intensity variation. Despite this, the qualitative results depicted in Figure 7, show that as in the case of sonar data, the masks are not ‘sharp’. It is noteworthy that the mask size is proportional to the object size. Notably, when employing AE-based clustering with a cluster number equal to 10, the dataset achieves the highest average precision at $83\%$, compared to $60\%$ when using true clusters.
To further discuss the reason for the superior performance for synthetic clusters, we show the clustered background using AE-K-means along with the corresponding true clusters in Figure 8. We notice that the K-means clustering creates clusters out of the same true cluster, and each cluster has a different phase. For example, in the case of 5 clusters, the brick cluster is divided into two out-of-phase clusters. It also groups together distinct true backgrounds into one cluster. Another example, in the case of 5 clusters, the first row, which represents a cluster, shows a mix of two true clusters. The sliced Wasserstein distance is insensitive to out-of-phase differences, making it difficult to quantify these differences. Consequently, images with identical content but differing phase are challenging to assign to the same distribution.

### v2

@@ Acknowledgments
Research at the University of Delaware was sponsored by the Department of the Navy, Office of Naval Research under ONR award numbers N00014-21-1-2300 and N00014-24-1-2259. This material is based upon work supported by the National Science Foundation under Award No. 2108841. This research was also supported in part through the use of DARWIN computing system: DARWIN – A Resource for Computational and Data-intensive Research at the University of Delaware and in the Delaware Region, which is supported by NSF under Grant Number: 1919839, Rudolf Eigenmann, Benjamin E. Bagozzi, Arthi Jayaraman, William Totten, and Cathy H. Wu, University of Delaware, 2021, URL.
@@ A.1 Generative-based Models
The work by Remez et al. (2018) is similar to ours in that it proposes a weakly-supervised approach that relies on a discriminator to assess the statistical similarity between synthetic images with counterfactual backgrounds and real images. Notably, their discriminator relies on a pretrained object classifier. Furthermore, the masking network input is based on the features from a pretrained object recognition network that outputs a bounding box for the foreground object. The necessity of using pretrained models and a network to generate a bounding box are the main limitations of this work.
@@
Bielski & Favaro (2019) proposed a GAN-based, unsupervised approach to segment foreground pixels. The GAN produces mask, foreground, and background images. The mask is used to alpha-blend the background and foreground back together. Trivial solutions for this work result when (i) the generated background and foreground are identical, so any arbitrary mask would still give a real image, (ii) Either the foreground or the background is ignored (i.e., all-foreground or all-background), which results in an either all-zero or all-one mask. To prevent the latter, the hinge loss on the average mask value is deployed. To prevent the former, a small random spatial shift of the composite image and mask are applied before alpha-blending to create a counterfactual shift in the resulting synthetic image. This shift penalizes having poor generated masks with identical foreground and background by creating a statistical divergence between shifted counterfactuals (i.e., the generated image) and real images, which the discriminator can detect. A drawback to this approach is the number of different loss functions (size, binary, gradient penalty, discriminator output norm) and hyperparameters required to prevent degenerate solutions.
@@
The work by Sauer & Geiger (2021) introduced a Counterfactual Generative Network (CGN), leveraging the concept of learning Independent Mechanisms (IMs) to establish a causal structure for mask, foreground, and background. Each was modeled by fine-tuning a distinct pretrained BigGAN. To ensure the resulting composite image is real, they minimize the MSE loss between the output of pretrained frozen BigGAN and the composition process. To ensure the generated mask is not trivial, hinge and entropy losses are deployed. For the background loss, the predicted saliency is minimized which leads to the foreground object’s disappearance. For background generation a pretrained U2Net is used to detect the saliency. Overall, this approach heavily relies on pretrained models and necessitates the incorporation of specific losses to prevent degenerate or trivial solutions.
@@ A.2 Vision-Transforms-based Models
The work by Hamilton et al. (2022) proposed an unsupervised semantic segmentation approach that distills knowledge from pretrained feature extractors. A frozen backbone (e.g., DINO-ViT) is used to extract dense pixel-wise features from input images. A trainable segmentation network maps these features to lower-dimensional embeddings representing per-pixel segment assignments. During training, a mini-batch is constructed for each sample by including its K-nearest neighbors in the feature space. The model then samples a neighbor image and computes a cosine similarity matrix between the pixel features of the target and the neighbor. The segmentation output is encouraged to be similar for pixels with high feature similarity and dissimilar for pixels with low similarity.
@@ A.3 Class-Activation-Map-based Models
Much work on weakly-supervised segmentation is based on class-activation maps (CAM) derived from pretrained convolutional object classification networks (Kumar Singh & Jae Lee, 2017; Wei et al., 2017; Zhang et al., 2018; Li et al., 2018; Hou et al., 2018; Jiang et al., 2019). We stress the fact that our model does not require labeled data (for instance, class information for composite/background images). Performance was initially limited as segmentation based on activation maps often focused on a few parts of the objects (Kumar Singh & Jae Lee, 2017; Wei et al., 2017). To mitigate this problem, Zhang et al. (2018) used a backbone fully-convolutional network to extract initial CAMs and then used two parallel-branches, each a CNN classification network, which took the initial CAM as an input. Parts of the input CAM of the first classifier are erased with the guidance of the object localization map from the second classifier. This encourages the first classifier to seek other parts of the image to do the classification. Thus, this results in learning a more coherent/complete feature map that can be used for object segmentation. Similarly, Li et al. (2018) proposed two networks, where one network uses the parts that are neglected by the first one for the classification task which results in expanding the activation map to other related parts. The work by Hou et al. (2018) introduced an approach for adversarial learning during supervised classification training which encourages the CAM to expand from covering only part of the target objects while penalizing it for covering background regions spuriously using the background prior (i.e., regions with low activations defined as background, regions with high activations defined as foreground and regions in-between defined as potential foreground) obtained from the initial map. Following a similar aim, Jiang et al. (2019) observed that the attention maps produced by a classification network changes focus on different object parts throughout the learning. They saved copies of the model each epoch and then used them to produce CAMs. At the end, they fused all the resulting CAMs using a pixel-wise max operation, which enhanced the final CAM. All previously mentioned work in this paragraph depends on pretrained models.
@@ A.4 Other Work
Hung et al. (2019) proposed SCOPS, a self-supervised part segmentation framework that leveraged multiple auxiliary objectives. First, they used saliency maps extracted from a pretrained classification network as soft supervision, encouraging the encoder’s part activation maps to align with semantically meaningful regions via a cosine similarity loss. Second, they introduced a geometric loss that encouraged all pixels assigned to a given part to be spatially concentrated around that part’s center; the number of parts was treated as a hyperparameter. Lastly, an equivariance loss ensured that part assignments remained consistent under random spatial transformations such as rotation and scaling, promoting robustness and part consistency.
The work by Savarese et al. (2021) proposed an unsupervised model-free method to segment foreground and background based on minimizing the mutual information between the segmented foreground and background. They used a variant of mutual information, namely the coefficient of constraint that prevents trivial solution including an entropy in the denominator. By including assumptions about the conditional distributions (i.e., Laplace distribution for foreground/background given background/foreground), the loss is equivalent to maximizing the inpainting errors of both the masked foreground and the masked background. However, they added a new loss that penalizes the texture diversity within the masked region (i.e., the masked region should have a uniform texture). The mask is estimated for each image by projected gradient descent on the mutual information loss.
@@ Appendix B Statistical Divergences
The family of Wasserstein distances is another class of divergences based on optimal transport theory (Villani, 2008). In particular, the Wasserstein-$1$ distance, a statistical divergence also known as the Earth mover’s distance, has been applied in generative modeling in the form of the Kantorovich dual (Arjovsky et al., 2017). Note that all of these divergences require adversarial optimization (MMD-GAN (Li et al., 2017) uses adversarial selection of the kernel). In contrast, the primal form of the Wasserstein-$p$ distance measures the divergence between the distributions of two random variables $X\sim\mathbb{P}_{X}$ and $Y\sim\mathbb{P}_{Y}$ in terms of the optimal transport plan
@@
$\overline{\mathrm{SW}}_{p}^{p}(\mathbb{P}_{X},\mathbb{P}_{Y})$ $\triangleq\int_{\mathbb{S}^{d-1}}\mathrm{SW}_{p}^{p}(\mathbb{P}_{X},\mathbb{P}_{Y};w)\,dw$ (11)
@@
Given two samples/batches $\hat{\mathbb{P}}_{X}=\{(\frac{1}{N},\mathbf{x}_{i})\}_{i=1}^{N}$ and $\hat{\mathbb{P}}_{Y}=\{(\frac{1}{N},\mathbf{y}_{i})\}_{i=1}^{N}$, the computation of the sliced Wasserstein-$p$ distance is simply the mean of the absolute error to the $p$th power between paired samples after sorting: $\mathrm{SW}_{p}^{p}(\hat{\mathbb{P}_{X}},\hat{\mathbb{P}}_{Y};\mathbf{w})=\frac{1}{N}\sum_{j=1}^{N}(\mathbf{w}^{\top}\mathbf{x}_{\pi_{j}}-\mathbf{w}^{\top}\mathbf{y}_{\sigma_{j}})^{p}$, where the permutations $\pi$ and $\sigma$ ensure that $\mathbf{w}^{\top}\mathbf{x}_{\pi_{1}}\leq\mathbf{w}^{\top}\mathbf{x}_{\pi_{2}}\leq\cdots\leq\mathbf{w}^{\top}\mathbf{x}_{\pi_{N}}$ and $\mathbf{w}^{\top}\mathbf{y}_{\sigma_{1}}\leq\mathbf{w}^{\top}\mathbf{y}_{\sigma_{2}}\leq\cdots\leq\mathbf{w}^{\top}\mathbf{y}_{\sigma_{N}}$, respectively. Intuitively, $\mathbf{w}$ defines a subspace and the sorting ensures the shortest distances between the pairs in the subspace.
Alternatively, one can use energy-based Wasserstein (EBSW) (Nguyen & Ho, 2023) that depends on importance sampling to emphasize slices that distinguish the distributions:
@@
$\sigma_{X,Y}(w;f,p)$ $\triangleq\frac{f(\text{SW}^{p}_{p}(\mathbb{P}_{X},\mathbb{P}_{Y};w))}{\int_{\mathbb{S}^{d-1}}f(\text{SW}^{p}_{p}(\mathbb{P}_{X},\mathbb{P}_{Y};w^{\prime}))dw^{\prime}},$ (13)
@@
Interestingly, an analogous approach can be used to select from a predefined set of kernel functions for MMD (Biggs et al., 2023), which provides a rough approximation of the adversarial optimization of the kernel function in MMD-GAN (Li et al., 2017).
@@ C.1 Synthetic Data Description
C.1.1 Synthetic-Aperture Sonar Backgrounds with dSprites (SAS+dSprites)
This synthetic dataset is derived from real-world background-only synthetic-aperture sonar (SAS) images (Cobb & Zare, 2014), augmented by overlaying artificial objects to simulate relevant scenarios. This SAS dataset contains 129 $1001\times 1001$ images that are originally complex-valued, single (high-frequency) channel, taken from various seafloor texture types. For each sonar scene, we crop 200 $64\times 64$ images from the magnitude images. We normalize each crop by its mean and then clip any value larger than 16. Half of the resulting crops are used as composite images where we paste objects from dSprites (Matthey et al., 2017) dataset using equation 1. The data is divided into $70\%,20\%,\text{ and }10\%$ for training, validation, and testing, respectively.
@@ C.1.1 Synthetic-Aperture Sonar Backgrounds with dSprites (SAS+dSprites)
C.1.2 Textures+MNIST/FashionMNIST/dSprites
@@
For the synthetic dataset, $140,000$ background images of size $64\times 64$ were generated by randomly patching the texture images. Half of the background images were used as background for the composite images. This resulted in $70,000$ images as foreground and background. Images in the foreground datasets were resized to $64\times 64$ and the true (binary) mask was used to blend it onto the background image using equation 1. Since MNIST and Fashion MNIST have no given true mask, we estimate the true mask by applying Otsu’s threshold on each image (Otsu, 1979). The data is divided into $70\%,20\%,\text{ and }10\%$ for training, validation, and testing, respectively.
@@
C.2 Implementation Details for Synthetic Datasets
@@
For SAS+dSprites and Textures+MNIST/FashionMNIST/dSprites, the surrogate data cluster labels for the composite images are computed by processing the corresponding background-only images through both the AE/CLE and Sinkhorn K-means. For the SAS images, we find that AE architecture is not able to reconstruct the images, so we use a symmetrical architecture that is composed of four blocks for the encoder and decoder, and each block is composed of convolution, ReLU, BatchNorm, with residual units: the 64-channel block has 1 and other blocks have 2 residual units each. The number of channels per block is $64,128,256,512$, with an additional 768-channel stage. As described above, a fully connected bottleneck layer of dimension $o=20$ is used to represent the latent space.
@@
The Textures+MNIST/FashionMNIST/dSprites images, have 10 ground truth clusters and we experiment with varying the number of clusters, choosing values from the set $\{5,10,15\}$. For SAS+dSprites images, we choose cluster values from the set $\{3,4,5\}$. Initial experiments with 2 clusters consistently failed to separate visually distinct regions.
@@
For Textures+MNIST/FashionMNIST/dSprites and SAS+dSprites, we run the training for $50$ and $100$ epochs, respectively. For the former, we choose the model with lowest validation loss $\mathcal{L}$ shown in equation 8, but for the latter we also include a quantile loss equation 15 with quantile, $q=0.8$, such that the maximum area of an object is 20% of the image size:
@@
$\mathcal{L}_{\text{quantile}}(\theta)$ $=\mathbb{E}_{X,R}\left[\sum_{ij\in\mathcal{Q}}\big([\tilde{X}_{\theta}]_{ij}-R_{ij}\big)^{2}\right],\quad\mathcal{Q}$ $=\left\{ij:M_{\theta}(X)_{ij}\leq\tau\right\},\quad\tau$ $=\text{quantile}\left(M_{\theta}(X),q\right).$ (15)
That is, for the SAS+dSprites dataset the total loss is $\mathcal{L}_{\mathit{D}}(\theta)+\mathcal{L}_{\text{Bg}}(\theta)+\mathcal{L}_{\text{quantile}}(\theta)$. This additional loss encourages the masking network to produce counterfactual images in which a $q$th fraction of its pixels match the source background, essentially encouraging the corresponding mask pixels to be zero.
C.3 Additional Results for Textures+MNIST/FashionMNIST/dSprites
Figure 10 shows the clustered background using AE+K-means along with the corresponding true clusters.